OpenStudy (stef1208):

Make sure you read the directions for each part carefully. This system of equations will be used in all three parts of this question. x-2y=0,5x+2y=24 Part 1: Solve the system by graphing. Draw the graphs on paper and indicate their intersection point. Type the ordered pair solution here as your answer to Part 1. Part 2: Solve the system using the substitution method. Show all work here and indicate the solution for the system as an ordered pair. Part 3: Solve the system using the addition method. Show all work here and indicate the solution for the system as an ordered pair

OpenStudy (rcp031):

Part 1: Solve the system by graphing. x-2y=0 ,5x+2y=24 Solve each equation for y y=x/2 & y=-5x/2+12 Identify slope and y −intercepts, Standard form y=mx+c m=slope=rise/run & c=constant m=1/2,c=0 & m=-5/2,c=12 y-intercepts To graph each equation, we start at the y-intercept and use the slope rise/run to get the next point and connect the dots.

OpenStudy (rcp031):

Part 2:The substitution method. x-2y=0,5x+2y=24 Find the lone variable y=x/2 Substitute into the untouched equation 5x+2*x/2=24 Solve for x 6x=24 x=4. Plug into lone variable equation and evaluate y=4/2=2. ordered pair (4,2)